A rectangle plot of ground measuring 12 meters by 20 meters is surrounded by a sidewalk of uniform width. The area of the sidewalk is 68 square meters. Find the width of the sidewalk
Answer by mananth(16946)

ground width = 12 m
length = 20 m
..
side walk width = xm
...
garden + sidewalk width = 12+2x m
garden+sidewalk length = 20+2x
..
Area of sidewalk = 68m^2
(12+2x)(20+2x)-20*12= 68
240+24x+40x+4x^2-240=68
4x^2+64x-68=0
/4
x^2+16x-17=0
x^2+17x-x-17=0
x(x+17)-1(x+17)=0
(x+17)(x-1)=0
x=-17 not possible
m.ananth@hotmail.ca
x= 1m the width of the sidewalk.